You are designing a miniature golf course and need to calculate the surface area and volume of many of the objects that will be used to build the course.
a. The first hole has a sphere that has a radius of 5 feet. What is the surface area and volume of the sphere?
b. The fourth hole has a pyramid that has a height of 12 feet; the base is a square with sides that measure 8 feet. What is the surface area and volume of the pyramid?
c. The seventh hole has a cone that has a height of 8 feet; the base has a radius of 5 feet. What is the surface area and volume of the cone?
d. The fourteenth hole has a rectangular solid that measures 10 feet long, 6 feet wide, and 16 feet tall. What is the surface are and volume of the rectangular solid?
e. Your boss wants to place signs on each face of the structures. How many faces are on the four geometric shapes on the holes?
f. To meet building codes we need to know the number of edges to place the appropriate brackets. How many edges will need brackets on the four shapes?
g. For extra support we also need to place additional brackets at each vertex. How many vertices are there for each structure?

Answer by addingup(3677)

a) Volume of a sphere:
(4Pi/3)r^3 = V
(4Pi/3)5^3 = 523.61
--------------------------
Surface of a sphere:
4Pi(r^2) = S
4Pi(5^2) = 314.16 (your teacher has a sense of humor. 3.1416 is the abbreviated value of Pi)
Sorry, it's too late and I need to get some sleep. Let me know if you don't get an answer to the rest of the problem and I will do it and send it to you.